OpenStudy (anonymous):

Explain, in complete sentences, how you would completely factor 20x2 – 28x – 48 by using the grouping method. How would you check your factors for accuracy?

OpenStudy (unklerhaukus):

take out a factor of four

OpenStudy (anonymous):

20x^2-28x-48*

OpenStudy (anonymous):

First of all, as you can see you can take out 4 as common from all the terms... Just do that..

OpenStudy (anonymous):

I alredy did @UR. It's just the next part which I'm stuck on. I got 4(x^2-7x-12).

OpenStudy (anonymous):

See, what is c here?? Compare it with the standard equation:

OpenStudy (anonymous):

4(5x^2-7x-12)

OpenStudy (anonymous):

Sry, that was a typo. Now that we have the GCF factored out, Idk what to do really.

OpenStudy (unklerhaukus):

now you can see 5 is a prime number so it must be a coefficient of one of the x's in the brackets in the next step

OpenStudy (unklerhaukus):

\[4(5x^2-7x-12)=4(5x+...)(x+...)\]

OpenStudy (anonymous):

ax^2 + bx^2 + c Tell me what is the value of c there??

OpenStudy (anonymous):

@water it's -12. and I see UR.

OpenStudy (anonymous):

No factors of -12 fulfill this.

OpenStudy (anonymous):

Oh really very sorry.. a is not 1 it is 5.. just multiply a and c : ac = 5*(-12) = -60 Now guess two numbers with product = -60 and sum = -7 See, I explain you: When a = 1, then we make the factors of c When a not equal to one then we make the factors of ac.. Getting this point?? just make the factors of -60 which should produce the sum = -7 and product = -60..

OpenStudy (anonymous):

I get it.

OpenStudy (anonymous):

As, this one is easy.. two numbers will be 12 and 5.. But we should need product = -60 When their product is -60?? When one of the number must be negative.. So either 12 can be negative or you can take 5 as negative.. But the other condition says that their sum should be -7.. With 12 and 5 how can you get sum = -7 It will only be possible if 12 is negative and 5 is positive so that -12 + 5 = -7 Got till here?